Describe how you would arrange a network of capacitors to store maximum charge.

(Use the Multiple Capacitors tab to design and test your hypothesis.)
Answer #1

We know that when two capacitors are connected in series its equivalent capacitance is given by

1/C =1/c1 + 1/c2

Where c1 and c2 are the capacitors. And when they are connected in parallel the equivalent capacitance is algebraic sum of the capacitors.

I.e C=c1+c2

Hence we conclude that when capacitors are connected in parallel the equivalent capacitance increases.

We know that the relationship between capacitance and charge is Q= C*V

Where Q=charge

C= Capacitance

V = voltage.

Hence we can see that charge is directly proportional to the capacitance if we keep the voltage constant hence the parallel connection of the capacitors will yield storing of more charge. And also in parallel connection voltage is higher as compared to the series connection. Hence we can assume any number of capacitors in parallel to store more charge with required values.

If your doubt is cleared please give an upvote it means a lot. Its a humble request to you.